# ORBGRAND Is Exactly Capacity-achieving via Rank Companding

### v2

@@ I Introduction
In this paper, we study a modification of ORBGRAND, which essentially adopts a companding technique to transform the rank of each LLR magnitude according to the inverse cumulative distribution function (CDF) of LLR magnitude. The resulting algorithm is thus termed CDF-ORBGRAND. The value of the shape of the CDF of LLR magnitude has been recognized earlier in [7], and a piecewise linear approximation technique has been utilized therein under the term of “full ORBGRAND”; see also [18, 19]. Via such a companding technique, the decoding metric is rendered asymptotically aligned with the channel law, while still maintaining its amenability to hardware implementation. We analyze the achievable rate of CDF-ORBGRAND for general binary-input memoryless channels under symmetric input distribution, and prove via a random coding argument akin to that in [11] that, CDF-ORBGRAND exactly achieves the mutual information, i.e., the symmetric capacity, without any rate loss. In contrast, for fading channels, ORBGRAND without rank companding may incur a significant rate loss compared to the symmetric capacity; see Fig. 1 in Section III-A.
On the basis of CDF-ORBGRAND for binary-input channels, we proceed to the investigation of bit-interleaved coded modulation (BICM) systems [20, 21]. This investigation is motivated by the prevalence of high-order modulation in modern communication systems, wherein BICM has been a widely used technique. By jointly considering the mismatched decoding effects due to both BICM and ORBGRAND, we derive achievable rates of both CDF-ORBGRAND and ORBGRAND for BICM. Remarkably, we show that CDF-ORBGRAND achieves the BICM capacity, which was initially derived in the literature by treating BICM as a set of independent parallel channels [20], and subsequently established by typicality decoding or by mismatched decoding with appropriately chosen metrics [22]. The BICM capacity-achieving property of CDF-ORBGRAND indicates that the rank companding technique is effective even for handling the heterogeneity among bit channels corresponding to different positions of the labeling of constellation symbols. In contrast, without rank companding, the rate loss of ORBGRAND is evident and becomes more substantial as modulation order increases; see Fig. 4 in Section IV-E.
@@ III-A Main Result
We conduct some numerical experiments to compare (4) and (5), for antipodal signaling, i.e., BPSK modulation, over AWGN, additive white Laplacian noise (AWLN), and Rayleigh fading channels with receiver-side perfect channel state information (CSI). These channels can be written as
@@
  • Rayleigh fading channel with receiver-side perfect CSI, $\mathsf{H}\sim\mathcal{CN}(0,1)$ and $\mathsf{Z}\sim\mathcal{CN}(0,1)$: $\mathsf{T}=4\sqrt{P}\Re(\mathsf{H}^{*}\mathsf{Y})$. We point out that here $q^{\pm}(\cdot)$ should be evaluated for the augmented channel output $(y,h)$, i.e., including the receiver-side perfect CSI. [^4] We clarify that in the numerical experiments in [25], the GMI of ORBGRAND for this setup was incorrectly computed, due to an error in the program evaluating the integrals in (4). In the current paper, we rectify the error and provide updated numerical results. The program code can be found at: https://github.com/ustclizhuang/Simulation
The curves of $I_{\text{ORB}}$ and $I_{\text{CDF-ORB}}$ for different signal-to-noise ratio (SNR) values are displayed in Fig. 1. Their comparison indicates that in AWGN channel, ORBGRAND almost achieves the symmetric capacity, a result already established in [11]; that in AWLN channel, a noticeable gap appears; and that for Rayleigh fading channel, the gap becomes even more pronounced.
@@ IV-E Numerical Results
The numerical results are shown in Fig. 4. It can be observed that ORBGRAND incurs a significant gap to the BICM capacity under both Gray and SP labelings. For example, with 8PSK and Gray labeling at SNR 15 dB, $I_{\text{ORB,BICM}}$ falls short of the BICM capacity (i.e., $I_{\text{CDF-ORB,BICM}}$) by approximately 0.5 bits per channel use. Moreover, this gap grows as the modulation order increases, reflecting the heightened sensitivity of ORBGRAND to high-order modulation in fading channels. Under the same conditions, the gap with SP labeling is slightly more pronounced than that with Gray labeling, indicating the impact of the choice of labeling scheme.
The performance gain of CDF-ORBGRAND over ORBGRAND can be attributed to the improved alignment of decoding metrics via rank companding, as commented at the beginning of Section III. This phenomenon appears to be the most evident in the intermediate regime of SNR, and as modulation order increases.
In Fig. 4 we also include the achievable rates $\tilde{I}_{\text{ORB,BICM}}$ of BICM derived in [25] which was derived under the assumption of ideal interleaving; [^5] Similar to Fig. 1, here we rectify an error in the numerical program in [25] and update the numerical results. see the end of Appendix E for details. Although the derivation relies upon an ideal assumption, we observe that the curves of $\tilde{I}_{\text{ORB,BICM}}$ are very close to the exact ones of ${I}_{\text{ORB,BICM}}$, indicating that the ideal interleaving assumption is usually sufficient for practical purposes.
@@ V Conclusion
In this work, we first analyze the achievable rates of ORBGRAND and its variant under rank companding, CDF-ORBGRAND, from the perspective of the GMI for general binary-input memoryless channels. A somewhat unexpected yet basic finding is that the achievable rate of CDF-ORBGRAND is identical to the symmetric capacity, indicating that rank companding loses no information asymptotically. Building upon this finding, we extend the analysis to BICM and demonstrate that, CDF-ORBGRAND also achieves the BICM capacity. By contrast, ORBGRAND sometimes experiences a substantial rate loss in fading channels.
@@ Appendix B Proofs of Lemmas 2, 5, and 8
For $i\neq j$, we have $\mathbb{E}[\tilde{\mathsf{W}}_{i}\tilde{\mathsf{W}}_{j}]=\mathbb{E}[\mathsf{W}_{i}\mathsf{W}_{j}]-\mathbb{E}[\mathsf{W}_{i}]\mathbb{E}[\mathsf{W}_{j}]$. We now turn to the evaluation of $\mathbb{E}[\mathsf{W}_{i}\mathsf{W}_{j}]$, which, by definition, can be expressed as (37), shown at the top of this page. The first term in (37) leads to (38) and (39), shown at the top of this page and the next page, respectively. Note that $\mathsf{R}_{i}$ and $\mathsf{R}_{j}$ in (39) are the ranks resulting from inserting $\lvert t_{i}\rvert=\bigl\lvert\ln\frac{q^{+}(y_{i})}{q^{-}(y_{i})}\bigr\rvert$ and $\lvert t_{j}\rvert=\bigl\lvert\ln\frac{q^{+}(y_{j})}{q^{-}(y_{j})}\bigr\rvert$ into a sorted array of N2N-2 i.i.d. samples of $\lvert\mathsf{T}\rvert$, and we denote by $u(y_{i},y_{j})$ the upper branch of (39). Therefore, (38) can be rewritten as $\int_{\mathcal{R}_{1}}\int_{\mathcal{R}_{1}}u(y_{i},y_{j})q^{+}(y_{i})\mathrm{d}y_{i}q^{+}(y_{j})\mathrm{d}y_{j}$. Applying a similar procedure to the other three terms in (37), we obtain

### v3

@@ I Introduction
In this paper, we study a modification of ORBGRAND, which essentially adopts a companding technique to transform the rank of each LLR magnitude according to the inverse cumulative distribution function (CDF) of LLR magnitude. The resulting algorithm is thus termed CDF-ORBGRAND. The value of the shape of the CDF of LLR magnitude has been recognized earlier in [7], and a piecewise linear approximation technique has been utilized therein under the term of “full ORBGRAND”; see also [18, 19]. Via such a companding technique, the decoding metric is rendered asymptotically aligned with the channel law, while still maintaining its amenability to hardware implementation. We analyze the achievable rate of CDF-ORBGRAND for general binary-input memoryless channels under symmetric input distribution, and prove via a random coding argument akin to that in [11] that, CDF-ORBGRAND exactly achieves the mutual information, i.e., the symmetric capacity, without any rate loss. Although numerical experiments suggest that for ORBGRAND, the rate gap to the symmetric capacity is already negligible for several representative channels, establishing the exactly capacity-achieving property for CDF-ORBGRAND is still reassuring from an information-theoretic perspective, shedding further insights into the behavior of ORBGRAND.
On the basis of CDF-ORBGRAND for binary-input channels, we proceed to the investigation of bit-interleaved coded modulation (BICM) systems [20, 21]. This investigation is motivated by the prevalence of high-order modulation in modern communication systems, wherein BICM has been a widely used technique. By jointly considering the mismatched decoding effects due to both BICM and ORBGRAND, we derive achievable rates of both CDF-ORBGRAND and ORBGRAND for BICM. Remarkably, we show that CDF-ORBGRAND achieves the BICM capacity, which was initially derived in the literature by treating BICM as a set of independent parallel channels [20], and subsequently established by typicality decoding or by mismatched decoding with appropriately chosen metrics [22]. The BICM capacity-achieving property of CDF-ORBGRAND indicates that the rank companding technique is effective even for handling the heterogeneity among bit channels corresponding to different positions of the labeling of constellation symbols.
@@ III-A Main Result
We conduct some numerical experiments to compare (4) and (5), for antipodal signaling, i.e., BPSK modulation, over AWGN, AWLN, and Rayleigh fading channels with receiver-side perfect channel state information (CSI). These channels can be written as
@@
  • Rayleigh fading channel with receiver-side perfect CSI, $\mathsf{H}\sim\mathcal{CN}(0,1)$ and $\mathsf{Z}\sim\mathcal{CN}(0,1)$: $\mathsf{T}=4\sqrt{P}\Re(\mathsf{H}^{*}\mathsf{Y})$. We point out that here $q^{\pm}(\cdot)$ should be evaluated for the augmented channel output $(y,h)$, i.e., including the receiver-side perfect CSI. [^4] The program code can be found at: https://github.com/ustclizhuang/Simulation
The curves of $I_{\text{ORB}}$ and $I_{\text{CDF-ORB}}$ for different signal-to-noise ratio (SNR) values are displayed in Fig. 1. Their comparison indicates that in AWGN and Rayleigh fading channels, the rate loss of ORBGRAND is essentially negligible, an observation already made in [11, 25], and that in AWLN channel, the gap is slightly widened.
@@ IV-E Numerical Results
The numerical results are shown in Fig. 4. It can be seen that, under both Gray and SP labelings, ORBGRAND exhibits only a slightly larger gap to the BICM capacity (i.e., $I_{\text{CDF-ORB,BICM}}$) in the low-SNR regime. Under the same conditions, the gap with SP labeling is slightly more pronounced than that with Gray labeling, indicating the impact of the choice of labeling scheme.
In Fig. 4 we also include the achievable rates $\tilde{I}_{\text{ORB,BICM}}$ of BICM derived in [25] which was derived under the assumption of ideal interleaving; see the end of Appendix E for details. Although the derivation relies upon an ideal assumption, we observe that the curves of $\tilde{I}_{\text{ORB,BICM}}$ are very close to the exact ones of ${I}_{\text{ORB,BICM}}$, indicating that the ideal interleaving assumption is usually sufficient for practical purposes.
@@ V Conclusion
In this work, we first analyze the achievable rates of ORBGRAND and its variant under rank companding, CDF-ORBGRAND, from the perspective of the GMI for general binary-input memoryless channels. A somewhat unexpected yet basic finding is that the achievable rate of CDF-ORBGRAND is identical to the symmetric capacity, indicating that rank companding loses no information asymptotically. Building upon this finding, we extend the analysis to BICM and demonstrate that, CDF-ORBGRAND also achieves the BICM capacity.
@@ Appendix B Proofs of Lemmas 2, 5, and 8
For $i\neq j$, we have $\mathbb{E}[\tilde{\mathsf{W}}_{i}\tilde{\mathsf{W}}_{j}]=\mathbb{E}[\mathsf{W}_{i}\mathsf{W}_{j}]-\mathbb{E}[\mathsf{W}_{i}]\mathbb{E}[\mathsf{W}_{j}]$. We now turn to the evaluation of $\mathbb{E}[\mathsf{W}_{i}\mathsf{W}_{j}]$, which, by definition, can be expressed as (37), shown at the top of this page. The first term in (37) leads to (38) and (39), shown at the top of this page. Note that $\mathsf{R}_{i}$ and $\mathsf{R}_{j}$ in (39) are the ranks resulting from inserting $\lvert t_{i}\rvert=\bigl\lvert\ln\frac{q^{+}(y_{i})}{q^{-}(y_{i})}\bigr\rvert$ and $\lvert t_{j}\rvert=\bigl\lvert\ln\frac{q^{+}(y_{j})}{q^{-}(y_{j})}\bigr\rvert$ into a sorted array of N2N-2 i.i.d. samples of $\lvert\mathsf{T}\rvert$, and we denote by $u(y_{i},y_{j})$ the upper branch of (39). Therefore, (38) can be rewritten as $\int_{\mathcal{R}_{1}}\int_{\mathcal{R}_{1}}u(y_{i},y_{j})q^{+}(y_{i})\mathrm{d}y_{i}q^{+}(y_{j})\mathrm{d}y_{j}$. Applying a similar procedure to the other three terms in (37), we obtain